OpenStudy (anonymous):

Express x = e^(-3t) , y= 5e^(2t) in the form y= f(x) by eliminating the parameter... need some help on simplifying for t on x = e^(-3t)...... Any help ????

OpenStudy (anonymous):

ln(x)=-3t t=ln(x)/-3 ln(y)=(ln(5)+2t) t=(ln(y)-ln(5))/2 t=t ln(x)/-3=(ln(y)-ln(5))/2 e^(ln(x)/-3)=e^((ln(y)-ln(5))/2) e^(ln(x)/-3)=sqrt(y/5) ...

OpenStudy (anonymous):

thanks for the reminder